Business Finance

Liz’s Health Food Store has estimated monthly financing requirements for the next six months as follows: 

 







  January$9,400    April$9,400  

  February3,400    May10,400  

  March4,400    June5,400  


 

Short-term financing will be utilized for the next six months. Projected annual interest rates are:

 







  January8.0%  April15.0%  

  February9.0%  May12.0%  

  March12.0%  June12.0%  


 

a. Compute total dollar interest payments for the six months. To convert an annual rate to a monthly rate, divide by 12. (Round intermediate calculations and final answers to 2 decimal places.)

 

Total dollar interest payments            $ 

 

b-1. Compute the total dollar interest payments if long-term financing at 12 percent had been utilized throughout the six months. Assume a long term rate is locked in on an interest-only loan.

 

Total dollar interest payments            $ 
Answer #1

Annual Amount Financed Monthly Month Monthly Rate Interest Rate Interest anuary February March April May 8% 9% 1290 15% 1290
Answer #2

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Σ Autobum a copy Formatting as Table Styles Styles Cells CN 740 WORKING CAPITAL APPROACH - Microsoft Excel (Product Activatio

Σ Autobum Copy 2 Clear Sort & Find & Filter Select Editing Cells CK CL CM | CN 740 WORKING CAPITAL APPROACH - Microsoft Excel
Answer #3

What is the answer for part B?
Answer #4

SOLUTION :


a.


Month  Financing required ($)    Interest rate           Interest amount ($)


Jan.               9400                                8%           9400*0.08/12     =  62.67


Feb.              3400                                 9%           3400*0.09/12     =  25.50


Mar.              4400                                12%          4400*0.12/12     =  44.00


Apr.              9400                                15%          9400*0.15/12     = 117.50


May             10400                               12%         10400*0.12/12    = 104.00


June             5400                                12%          5400*0.12/12     =  54.00


                                             Total interest amount                       = 407.67


Total interest amount for 6 months = $407.67 (ANSWER).


b1.


Total finance required in 6 months :    $42400 ($ * months)

Interest rate per month = 12%/12 = 1% = 0.01  ($ / ($.month))


Interest amount on longterm basis = 42400*0.01 =  424.00 ($) 


So, long-term basis interest amount  for 6 months = $424.00 (ANSWER).


answered by: Tulsiram Garg